OpenStudy (anonymous):

y=2x+5 Find the slope

OpenStudy (anonymous):

@mathmale @bloopman

OpenStudy (anonymous):

Too raw. Have you tried taking the line integral over [0, 2pi] for both sides?

OpenStudy (anonymous):

I used u-substitution to find the initial state of a qubit before it is observed in the plus/minus basis and used Grover's algorithm to search through a lattice of N indexes. But all that did was spawn a super massive black hole with me in the center of it. Now I'm stuck trying to find my way to the event horizon.

OpenStudy (anonymous):

I used to work for NASA until a few mathematical errors with the Challenger...

OpenStudy (anonymous):

I have no idea why though my logic is perfectly sane

OpenStudy (anonymous):

@tHe_FiZiCx99

OpenStudy (anonymous):

First, you'll probably need to verify that the integral is defined over the real plane. Then, let f(x,y) be a real-valued continuous function on the rectangle R = {(x,y); |x-x0| <= a; |y-y0| <= b}. Now, can you assume f has a partial derivative with respect to x?

OpenStudy (anonymous):

Also, have you tried turning down?

OpenStudy (anonymous):

I've tried... but my son keeps telling me turn down for what

OpenStudy (anonymous):

The slope is 2x. (0, 5) is the y intercept. Slope is normally described by the ratio of the rise divided by the run between two points on a line.

OpenStudy (the_fizicx99):

._. Luigi why didn't you answer X_X y = mx + b M being the slope b is the y intercept

OpenStudy (anonymous):

THANK YOU! Now I can escape this black hole via Hawking Radiation!

OpenStudy (the_fizicx99):

^

OpenStudy (anonymous):

Anddddd I'm in another black hole

OpenStudy (anonymous):

Good Luck!

OpenStudy (anonymous):

That's my son!